The monthly payments on a 10-year loan of $25,000 at 5.2% interest are $267.61.
(a) What is the total amount paid over the 10 years?
$
(b) What is the total amount of interest paid?
$

Answer by ikleyn(52879)

.
The monthly payments on a 10-year loan of $25,000 at 5.2% interest are $267.61.
(a) What is the total amount paid over the 10 years?
(b) What is the total amount of interest paid?
~~~~~~~~~~~~~~~~


        It is an  ENTERTAINMENT  type problem :
        to solve it,  you should not make complicated computations.


(a)  To answer (a), simply multiply $267.61 by 10*12 months to get

          Paid total = 10*12*267.61 = 32113.20 dollars.



(b)  To answer (b), simply subtract $25,000 from 32113.20 dollars to get

          total amount of interest paid = 32113.20 - 25000 = 7113.20 dollars.


Solved,  answered and explained with minimum calculations.

Have fun and smile.   Nice joke problem.